# Conformal Prediction for Ensembles: Improving Efficiency via Score-Based Aggregation

### v2

@@ Abstract
Distribution-free uncertainty estimation for ensemble methods is increasingly desirable due to the widening deployment of multi-modal black-box predictive models. Conformal prediction is one approach that avoids such distributional assumptions. Methods for conformal aggregation have in turn been proposed for ensembled prediction, where the prediction regions of individual models are merged as to retain coverage guarantees while minimizing conservatism. Merging the prediction regions directly, however, sacrifices structures present in the conformal scores that can further reduce conservatism. We, therefore, propose a novel framework that extends the standard scalar formulation of a score function to a multivariate score that produces more efficient prediction regions. We then demonstrate that such a framework can be efficiently leveraged in both classification and predict-then-optimize regression settings downstream and empirically show the advantage over alternate conformal aggregation methods.
@@ 1 Introduction
One method for performing distribution-free uncertainty quantification is conformal prediction, which provides a principled framework for producing distribution-free prediction regions with marginal frequentist coverage guarantees [23, 24]. By using conformal prediction on a user-defined score function, prediction regions attain marginal coverage guarantees. While calibration is guaranteed from this procedure, predictive efficiency, quantified as the size of the resulting prediction regions, can be unboundedly large for poorly chosen score functions.
@@ 2.1 Conformal Prediction
Coverage guarantees of uncertainty quantification methods generally rely on distributional assumptions, often via asymptotics or explicit specification. To alleviate the need for such restrictive assumptions, interest in finite-sample, distribution-free uncertainty quantification methods has risen. Conformal prediction is one such method [23, 24]. Due to its lack of structural assumptions, conformal prediction pairs well with the increasingly common black-box style of modeling.
@@ 2.2 Quantile Envelopes
Generalizations of quantiles have a long history in statistics [30, 31]. Unlike univariate data, multivariate data does not lend itself to an unambiguous definition of a quantile, as there is no canonical ordering in higher dimensional spaces. To account for this, one class of multivariate quantiles, known as “direction quantiles,” prescribes directions along which data points can be compared [32, 33, 34]. That is, points $x\in\mathbb{R}^{n}$ are projected along some direction $u\in\mathcal{S}^{n-1}$, after which the standard notion of a quantile can be leveraged.
The notion of the $\alpha$ quantile for a random variable $X\in\mathcal{X}$ clearly then depends on this choice of direction $u$, which we denote by $Q(X,\alpha,u)=\inf\{q\in\mathbb{R}:\mathcal{P}(u^{\top}X\leq q)\geq\alpha\}$. When there is no ambiguity, we just denote it as $Q(\alpha,u)$. In cases where domain knowledge assigns an unambiguous notion of multivariate ordering, such knowledge can be encoded into the choice of $u$. In other cases, however, a more agnostic approach is employed: rather than choosing any particular direction, all directions are considered simultaneously. For any given $u$, notice the choice of quantile defines a corresponding halfplane $H(u,Q(\alpha,u))=\left\{x\in\mathcal{X}:u^{\top}x\leq Q(\alpha,u)\right\}$. To consider all directions simultaneously, the quantile envelope is then defined to be
@@ 2.3 Predict-Then-Optimize
Prediction regions produced from conformalized predictors are often leveraged differently across classification and regression settings. In the case of classification, “regions” simply constitute a subset of the label space, making their direct use by end users straightforward, as demonstrated in works such as [35]. In high-dimensional regression settings, however, prediction regions become harder to use directly; for this reason, recent works have started shifting focus to using them in their implicit forms.
One such application is [36], where conformal prediction was leveraged in a predict-then-optimize setting. As the name suggests, predict-then-optimize problems are two-stage problems, which take observed contextual information $x$ and predict the parametric specification of a downstream problem of interest $\widehat{c}:=g(x)$ with some trained predictor $g$. The final result is then a decision made with this specification, $w^{*}:=\min_{w}f(w,\widehat{c})$. An example of such a setting is if an optimal labor allocation $w^{*}$ is sought based on predicted demand $\widehat{c}$ from observed transactions $x$ in a food delivery platform.
While the predicted $\widehat{c}$ is often simply trusted in the downstream optimization, this approach is inappropriate in risk-sensitive settings, where misspecification of the map $g:\mathcal{X}\rightarrow\mathcal{C}$ could lead to suboptimal decision-making. For this reason, recent interest has been placed on studying a “robust” alternative to this formulation [37, 38, 39]. Following this line of work, [36] proposed studying $w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, with $\mathcal{C}(x)$ being produced by conformalizing the predictor $g$. The primary consequence of doing so was having a probabilistic guarantee on the suboptimality gap. That is, denoting the suboptimality gap by $\Delta(x,c):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})-\min_{w}f(w,c)$, where $c$ is the true parameter corresponding to $x$, they established that leveraging conformal uncertainty regions guarantees $\mathcal{P}_{X,C}\left(0\leq\Delta(X,C)\leq L\mathrm{\ diam}(\mathcal{C}(X))\right)\geq 1-\alpha$ if $f(w,c)$ is an objective function that is convex-concave and $L$-Lipschitz in $c$ for any fixed $w$.
@@ 2.4 Related Works
Ensemble methods consist of $K$ predictors $f_{k}:\mathcal{X}_{k}\rightarrow\mathcal{Y}$; notably, such predictors need not map from the same set of covariates. A naive approach for uncertainty quantification would then be to conformalize the ensembled predictor. That is, for an ensembling algorithm $\mathcal{F}:\mathcal{Y}^{K}\rightarrow\mathcal{Y}$, a score function $s(\mathcal{F}(f_{1}(x),...,f_{K}(x)),y)$ would be defined. Denoting the $\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil/N_{\mathcal{C}}$ quantile of the score distribution over $\mathcal{D}_{C}$ as $\widehat{q}(\alpha)$, $\mathcal{C}(x)=\{y:s(x,y)\leq\widehat{q}(\alpha)\}$ would then be marginally calibrated.
@@
Conformal model aggregation, thus, seeks to mitigate these deficiencies by aggregating the prediction regions $\mathcal{C}_{1}(x),...,\mathcal{C}_{K}(x)$ rather than the individual point predictions [25, 27, 28, 29]. While there are several methods in this vein, they can be categorized into one of two general approaches. The first line of work seeks to perform model selection, in which a single conformal predictor is selected $\mathcal{C}_{k^{*}}$, typically based on the criterion of minimizing region size $k^{*}:=\operatorname*{arg\,min}_{k}\mathbb{E}[\mathcal{L}(\mathcal{C}_{k}(X))]$ [27, 28]. Generally, however, methods leveraging the full collection of predictors produce less conservative regions [25, 29]. Such works aggregate the individual prediction regions into a final region by defining
for weights $\{w_{k}\}\in[0,1]$ such that $\sum_{k=1}^{K}w_{k}=1$ and a threshold $\widehat{a}$. Methods then differ in the procedure by which $\{w_{k}\}$ and $\widehat{a}$ are prescribed, several of which were prescribed in [29], whose detailed presentation is deferred to Appendix F for space reasons. We note that the methods of [25] are designed for a different setting than that considered herein, namely that in which conformal coverage is sought adaptively over online data streams.
In this vein, [40] is a recent work that has similarly proposed a vector-score extension as that discussed herein, in which candidate weight vectors $\{w_{m}\}\in\mathbb{R}^{K}$ are searched over for score aggregation. That is, a vector $s(x):=(s_{1}(x,y),...,s_{K}(x,y))\in\mathbb{R}^{K}$ of scores $s_{k}(x,y)$ corresponding to each predictor $f_{k}(x)$ is predicted and its aggregate prediction region defined on the projection $\langle w_{m^{*}},s\rangle$ for $w_{m^{*}}$ the weight resulting in the smallest prediction region. This method, however, has two shortcomings addressed herein. The first is that their method can only be applied in classification settings, whereas our method can be leveraged across both regression and classification problems. The second is that their approach only uses a single weighted projection in the end, resulting in suboptimal aggregation and, therefore, conservative prediction regions.
@@ 3 Method
We now introduce a procedure for conformal score aggregation for ensembles. We discuss the approach and its coverage guarantees in Section 3.1. We then discuss the downstream use of its prediction regions in Section 3.2.
@@ 3.1 Multivariate Score Quantile
We wish to consider the setting typical of conformal model aggregation, as discussed in Section 2.4, in which predictors $f_{1}(x),...,f_{K}(x)$ and corresponding scores $s_{1}(x,y),...,s_{K}(x,y)$ are defined. We further assume a similar premise as that taken by [40], in which the individual scores are stacked into a multivariate score $s(x,y):=(s_{1}(x,y),...,s_{K}(x,y))$. A naive approach would then involve leveraging standard conformal prediction over a pre-defined map $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$, e.g., $g(s)=\sum_{k=1}^{K}s_{k}$. Similar to the naive conformalization of an ensembled predictor discussed in Section 2.4, using a fixed $g$ fails to adapt to any disparities in uncertainties present across predictors or requires intimate knowledge of such uncertainties. We instead wish to provide a data-adaptive pipeline to automatically produce such a $g$.
@@ 3.1.1 Score Partial Ordering
Intuitively, our method seeks to directly generalize the approach of split conformal, by “ordering” the collection of multivariate calibration scores and taking the $1-\alpha$ score under such an ordering to be a threshold $\widehat{\mathcal{Q}}$ with which prediction regions are then implicitly defined. Formally, the multivariate “ordering” is established as a pre-ordering $\lesssim$ over $\mathbb{R}^{K}$; a pre-ordering differs from a total ordering in that it need not satisfy the antisymmetric axiom of a total ordering. Roughly speaking, an “acceptance region,” so called as it serves as the criterion used to ultimately decide which $y$ are accepted into the prediction region, is then defined as $\widehat{\mathcal{Q}}:=\{s\mid s\lesssim\widehat{q}\}$, where $\widehat{q}$ is the $1-\alpha$ empirical quantile of $\mathcal{S}_{C}$ under $\lesssim$. Such a $\widehat{\mathcal{Q}}$ naturally generalizes the standard scalar acceptance interval of $[0,\widehat{q}]$ in the case of non-negative score functions. We briefly highlight the distinction between acceptance regions and prediction regions: the former are subsets of the space of multivariate scores, i.e. $\subset\mathbb{R}^{K}$, that ultimately define the criteria for retaining particular $y$ values in the prediction region whereas the latter are the subsets of the output space that ultimately have coverage guarantees, i.e. $\subset\mathcal{Y}$. The two, however, are directly related; in particular, for a fixed score $s(x,y)$, a larger acceptance region will result in a more conservative prediction region.
@@
For a family of sets $\{\mathcal{A}_{t}\}_{t\in\mathbb{R}}$, we take each $\mathcal{A}_{t}$ to be the region of the positive orthant $\mathbb{R}_{+}^{K}$ bounded by the coordinate axes and an “outer frontier” parameterized by $t$. The shape of this outer frontier remains fixed over the family and is merely scaled outward from the origin with $t$. Under this choice, performing a comparison of $s_{1},s_{2}\in\mathbb{R}^{K}$ in the aforementioned fashion, i.e. checking if $s_{1}\lesssim s_{2}$, amounts to checking if $t(s_{1})\leq t(s_{2})$, where $t(s)$ is the smallest $t$ for which the outer frontier of $\mathcal{A}_{t}$ intersects $s$. Notably, $t(s)$ is precisely the aforementioned data-driven score fusion function $g(s)$ of interest. Defining a data-adaptive $g(s)$, therefore, reduces to having a data-driven approach for defining the outer frontier of $\mathcal{A}_{t}$. We restrict this outer frontier to be such that $\mathcal{A}_{t}$ is a convex set; if $\mathcal{A}_{t}$ were permitted to be nonconvex, computing $t(s):=\min\{t\in\mathbb{R}:s\in\mathcal{A}_{t}\}$ would potentially be computationally expensive. The benefits of such convexity are highlighted, for example, in Section 3.2.
@@
This final reframing, therefore, naturally motivates selecting the outer frontier to be the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$. Using $\mathcal{S}_{C}$ to define $\mathcal{A}_{t}$ and in turn $\lesssim$, however, sacrifices the exchangeability of its points with future test scores $s^{\prime}$, as the very nature of ordering would change in swapping $s^{\prime}$ with any $s\in\mathcal{S}_{C}$. The goal, then, follows as seeking to define the outer frontier as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$ without directly using $\mathcal{S}_{C}$. For this reason, we partition $\mathcal{S}_{C}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$, where we define $\lesssim$ using $\mathcal{S}_{C}^{(1)}$ and compute the final $\widehat{q}$ over $\mathcal{S}_{C}^{(2)}$. Such a split is predicated on the assumption that the $1-\alpha$ quantile envelope defined over $\mathcal{S}_{C}^{(1)}$ resembles that of $\mathcal{S}_{C}^{(2)}$, implying the $|\mathcal{S}_{C}^{(1)}|$ should be sufficiently large as to capture this structure accurately.
We now focus attention on defining the quantile envelope over $\mathcal{S}_{C}^{(1)}$ using a technique paralleling that described in Section 2.2. In particular, we start by selecting the projection directions $\{u_{m}\}$ of Equation 1. The restriction of $s$ to the positive orthant induces a natural modification of Equation 1, namely where $u\in\mathcal{S}_{+}^{K-1}:=\mathcal{S}^{K-1}\cap\mathbb{R}_{+}^{K}$ instead of $\mathcal{S}^{K-1}$. To best approximate Equation 1, we wish for $\{u_{m}\}$ to be uniformly distributed over $\mathcal{S}_{+}^{K-1}$; however, exactly finding an evenly distributed set of points over hyperspheres in arbitrary $n$-dimensional spaces is a classically difficult problem [41] If $K=2$, we can solve this exactly; for $K>2$, we generate directions stochastically such that $U\sim\mathrm{Unif}(\mathcal{S}_{+}^{K-1})$ by drawing $V_{1},...,V_{M}\sim\mathcal{N}(0,I^{K\times K})$ and defining $U_{i}:=V_{i}^{|\cdot|}/\sqrt{V_{1}^{2}+...+V_{M}^{2}}$, where $v^{|\cdot|}$ denotes the component-wise absolute values of $v$.

### v3

@@ Abstract
Distribution-free uncertainty estimation for ensemble methods is increasingly desirable due to the widening deployment of multi-modal black-box predictive models. Conformal prediction is one approach that avoids such distributional assumptions. Methods for conformal aggregation have in turn been proposed for ensembled prediction, where the prediction regions of individual models are merged as to retain coverage guarantees while minimizing conservatism. Merging the prediction regions directly, however, can miss out on opportunities to further reduce conservatism by exploiting structures present in the conformal scores. We, therefore, propose a novel framework that extends the standard scalar formulation of a score function to a multivariate score that produces more efficient prediction regions. We then demonstrate that such a framework can be efficiently leveraged in both classification and predict-then-optimize regression settings downstream and empirically show the advantage over alternate conformal aggregation methods.
@@ 1 Introduction
One method for performing distribution-free uncertainty quantification is conformal prediction, which provides a principled framework for producing distribution-free prediction regions with marginal frequentist coverage guarantees [23, 24]. By using conformal prediction on a user-defined score function, prediction regions attain marginal coverage guarantees. While calibration is guaranteed from this procedure, predictive efficiency, i.e. the size of the resulting prediction regions, can be unboundedly large for poorly chosen score functions.
@@ 2.1 Conformal Prediction
Coverage guarantees of uncertainty quantification methods generally rely on distributional assumptions, often via asymptotics or explicit specification. To alleviate the need for such restrictive assumptions, interest in finite-sample, distribution-free uncertainty quantification methods has risen. Conformal prediction is one such method [23, 24].
@@ 2.2 Quantile Envelopes
Generalizations of quantiles have a long history in statistics [30, 31]. Unlike univariate data, multivariate data do not lend itself to an unambiguous definition of a quantile, as there is no canonical ordering in higher dimensional spaces. The notion of a “directional quantile” for a random variable $X\in\mathbb{R}^{n}$ can, however, be directly defined given some direction $u\in\mathcal{S}^{n-1}$, namely as $Q(X,\alpha,u)=\inf\{q\in\mathbb{R}:\mathcal{P}(u^{\top}X\leq q)\geq\alpha\}$ [32, 33, 34]. When there is no ambiguity, we just denote it as $Q(\alpha,u)$. For any given $u$, notice the choice of quantile defines a corresponding halfplane $H(u,Q(\alpha,u))=\left\{x\in\mathcal{X}:u^{\top}x\leq Q(\alpha,u)\right\}$. The quantile envelope is then the intersection thereof:
@@ 2.3 Predict-Then-Optimize
In the case of classification, conformal prediction regions simply constitute a subset of the label space, making their direct use by end users straightforward [35]. In high-dimensional regression settings, however, prediction regions become harder to use directly; for this reason, recent works have started shifting focus to using them in their implicit forms.
One such application is [36], where conformal prediction was leveraged in a predict-then-optimize setting. As the name suggests, predict-then-optimize problems are two-stage problems, which take observed contextual information $x$ and predict the parametric specification of a downstream problem of interest $\widehat{c}:=g(x)$ with some trained predictor $g$. The final result is then a decision made with this specification, $w^{*}:=\min_{w}f(w,\widehat{c})$. An example of such a setting is if an optimal labor allocation $w^{*}$ is sought based on predicted demand $\widehat{c}$ from transactions $x$ in a delivery platform.
While the predicted $\widehat{c}$ is often trusted, this approach is inappropriate in risk-sensitive settings, where misspecification of the map $g:\mathcal{X}\rightarrow\mathcal{C}$ could lead to suboptimal decision-making. For this reason, recent interest has been placed on studying a “robust” formulation [37, 38, 39]. Following this line of work, [36] proposed studying $w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, with $\mathcal{C}(x)$ being produced by conformalizing the predictor $g$.
@@ 2.4 Related Works
Ensemble methods consist of $K$ predictors $f_{k}:\mathcal{X}_{k}\rightarrow\mathcal{Y}$; notably, such predictors need not map from the same set of covariates. A naive approach for uncertainty quantification would then be to conformalize the ensembled predictor. That is, for an ensembling algorithm $\mathcal{F}:\mathcal{Y}^{K}\rightarrow\mathcal{Y}$, a score function $s(\mathcal{F}(f_{1}(x),...,f_{K}(x)),y)$ would be defined. Denoting the $\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil/N_{\mathcal{C}}$ quantile of the score distribution over $\mathcal{D}_{C}$ as $\widehat{q}(\alpha)$, $\mathcal{C}(x)=\{y:s(x,y)\leq\widehat{q}(\alpha)\}$ would then be calibrated.
@@
Conformal model aggregation, thus, seeks to mitigate these deficiencies by aggregating the prediction regions $\mathcal{C}_{1}(x),...,\mathcal{C}_{K}(x)$ rather than the individual point predictions [25, 27, 28, 29]. While there are several methods in this vein, they can be categorized into one of two general approaches. The first line of work seeks to perform model selection, in which a single conformal predictor is selected $\mathcal{C}_{k^{*}}$, typically based on the criterion of minimizing region size $k^{*}:=\operatorname*{arg\,min}_{k}\mathbb{E}[\mathcal{L}(\mathcal{C}_{k}(X))]$ [27, 28].
Generally, however, methods leveraging the full collection of predictors produce less conservative regions [25, 29]. Such works aggregate the individual prediction regions into a final region by defining $\mathcal{C}(x):=\{y\mid\sum_{k=1}^{K}w_{k}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]\geq\widehat{a}\}$ for weights $\{w_{k}\}\in[0,1]$ such that $\sum_{k=1}^{K}w_{k}=1$ and a threshold $\widehat{a}$. Methods then differ in the procedure by which $\{w_{k}\}$ and $\widehat{a}$ are prescribed, several of which were prescribed by [29], whose detailed presentation is deferred to Appendix L for space reasons. We note that the methods of [25] are designed for a different setting than that considered herein, namely that in which conformal coverage is sought adaptively over data streams.
In this vein, [40] have recently proposed a vector-score extension as that discussed herein, in which candidate weight vectors $\{w_{m}\}\in\mathbb{R}^{K}$ are searched over for score aggregation. That is, a vector $s(x):=(s_{1}(x,y),...,s_{K}(x,y))\in\mathbb{R}^{K}$ of scores $s_{k}(x,y)$ corresponding to each predictor $f_{k}(x)$ is predicted and its aggregate prediction region defined on the projection $\langle w_{m^{*}},s\rangle$ for $w_{m^{*}}$ the weight resulting in the smallest prediction region. This method, however, has two shortcomings addressed herein. The first is that their method can only be applied in classification settings, whereas our method can be leveraged across both regression and classification problems. The second is that their approach only uses a single weighted projection in the end, resulting in suboptimal aggregation and, therefore, conservative prediction regions.
@@ 3.1 Multivariate Score Quantile
We consider the setting typical of conformal model aggregation, as discussed in Section 2.4, in which predictors $f_{1}(x),...,f_{K}(x)$ and corresponding scores $s_{1}(x,y),...,s_{K}(x,y)$ are defined. We assume a similar premise as [40], in which the scores are stacked into a multivariate score $s(x,y):=(s_{1}(x,y),...,s_{K}(x,y))$. A naive approach would then leverage standard conformal prediction over a pre-defined map $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$, e.g., $g(s)=\sum_{k=1}^{K}s_{k}$. Similar to the naive conformalization of an ensembled predictor discussed in Section 2.4, using a fixed $g$ fails to adapt to any disparities in uncertainties present across predictors or requires intimate knowledge of such uncertainties. We instead wish to provide a data-adaptive pipeline to automatically produce such a $g$.
@@ 3.1.1 Score Partial Ordering
Intuitively, our method seeks to directly generalize the approach of split conformal, by “ordering” the collection of multivariate calibration scores and taking the $1-\alpha$ score under such an ordering to be a threshold $\widehat{\mathcal{Q}}$ with which prediction regions are then implicitly defined. Formally, the multivariate “ordering” is established as a pre-ordering $\lesssim$ over $\mathbb{R}^{K}$; a pre-ordering differs from a total ordering in that it need not satisfy the antisymmetric axiom of a total ordering. Roughly speaking, an “acceptance region,” so called as it serves as the criterion used to ultimately decide which $y$ are accepted into the prediction region, is then defined as $\widehat{\mathcal{Q}}:=\{s\mid s\lesssim\widehat{q}\}$, where $\widehat{q}$ is the $1-\alpha$ empirical quantile of $\mathcal{S}_{C}$ under $\lesssim$. Such a $\widehat{\mathcal{Q}}$ naturally generalizes the standard scalar acceptance interval of $[0,\widehat{q}]$ in the case of non-negative score functions. We briefly highlight the distinction between acceptance regions and prediction regions. The former are subsets of the space $\subset\mathbb{R}^{K}$ of multivariate scores that ultimately define the criteria for retaining particular $y$ values in the prediction region. The latter are the subsets of the output space $\mathcal{Y}$ and it is these that ultimately have coverage guarantees. The two, however, are directly related; in particular, for a fixed score $s(x,y)$, a larger acceptance region will result in a more conservative prediction region.
@@
For a family of sets $\{\mathcal{A}_{t}\}_{t\in\mathbb{R}}$, we take each $\mathcal{A}_{t}$ to be the region of the positive orthant $\mathbb{R}_{+}^{K}$ bounded by the coordinate axes and an “outer frontier” parameterized by $t$. The shape of this outer frontier remains fixed over the family and is merely scaled outward from the origin with $t$. Under this choice, comparing $s_{1},s_{2}\in\mathbb{R}^{K}$, i.e. checking if $s_{1}\lesssim s_{2}$, amounts to checking if $t(s_{1})\leq t(s_{2})$, where $t(s)$ is the smallest $t$ for which the outer frontier of $\mathcal{A}_{t}$ intersects $s$. Notably, $t(s)$ is precisely the aforementioned data-driven score fusion function $g(s)$ of interest. Defining a data-adaptive $g(s)$, therefore, reduces to having a data-driven approach for defining the outer frontier of $\mathcal{A}_{t}$. We restrict this outer frontier to be such that $\mathcal{A}_{t}$ is a convex set; if $\mathcal{A}_{t}$ were permitted to be nonconvex, computing $t(s):=\min\{t\in\mathbb{R}:s\in\mathcal{A}_{t}\}$ would potentially be computationally expensive. The benefits of such convexity are highlighted, for example, in Section 3.2.
@@
This motivates selecting the outer frontier to be the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$. Using $\mathcal{S}_{C}$ to define $\mathcal{A}_{t}$ and in turn $\lesssim$, however, sacrifices the exchangeability of its points with test scores $s^{\prime}$, as the very nature of ordering would change in swapping $s^{\prime}$ with any $s\in\mathcal{S}_{C}$. The goal follows as seeking to define the outer frontier as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$ without directly using $\mathcal{S}_{C}$. For this reason, we partition $\mathcal{S}_{C}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$, where we define $\lesssim$ using $\mathcal{S}_{C}^{(1)}$ and compute $\widehat{q}$ over $\mathcal{S}_{C}^{(2)}$. Such a split is predicated on the assumption that the $1-\alpha$ quantile envelope defined over $\mathcal{S}_{C}^{(1)}$ resembles that of $\mathcal{S}_{C}^{(2)}$, implying the $|\mathcal{S}_{C}^{(1)}|$ should be sufficiently large as to capture this structure accurately.
We now focus attention on defining the quantile envelope over $\mathcal{S}_{C}^{(1)}$ using a technique paralleling that described in Section 2.2. In particular, we start by selecting the projection directions $\{u_{m}\}$ of Equation 1; since $s\in\mathbb{R}_{+}^{K}$, we similarly restrict $u_{m}\in\mathcal{S}_{+}^{K-1}:=\mathcal{S}^{K-1}\cap\mathbb{R}_{+}^{K}$. To best approximate Equation 1, we wish for $\{u_{m}\}$ to be uniformly distributed over $\mathcal{S}_{+}^{K-1}$; however, exactly finding an evenly distributed set of points over hyperspheres in arbitrary $n$-dimensional spaces is a classically difficult problem [41] If $K=2$, we can solve this exactly; for $K>2$, we generate directions stochastically such that $U\sim\mathrm{Unif}(\mathcal{S}_{+}^{K-1})$ by drawing $V_{1},...,V_{M}\sim\mathcal{N}(0,I^{K\times K})$ and defining $U_{i}:=V_{i}^{|\cdot|}/\sqrt{V_{1}^{2}+...+V_{M}^{2}}$, where $v^{|\cdot|}$ denotes the component-wise absolute values.